OpenStudy (littlenugget):

In exercise 23, find the real numbers x and y that make the equation true: (5-2i)-7=x-(3+yi)

OpenStudy (anonymous):

Regrouping things: 5-7+3-2i=x-iy--->x=5-7+3 and y=2

OpenStudy (littlenugget):

Can you explain again? sorry i got a little confused :3 @CarlosGP

OpenStudy (anonymous):

Sure. This is an equation that is using complex numbers,agree?

OpenStudy (littlenugget):

yep

OpenStudy (anonymous):

and you know that complex numbers have two parts: Real part and Imaginary part, agree?

OpenStudy (littlenugget):

yep

OpenStudy (anonymous):

How do you identify the imaginary part?

OpenStudy (littlenugget):

with i?

OpenStudy (anonymous):

Correct! Then they are in fact giving you two equations: One for the imaginary part (all the terms multiplied by i) and another for the rest (the real part). Does it make sense to you?

OpenStudy (littlenugget):

i think, so in this equation, which would be the imaginary part and which would be the rest?

OpenStudy (anonymous):

The imaginary part are all the terms multiplied by "i". Can you identify them?

OpenStudy (littlenugget):

2i, and yi

OpenStudy (anonymous):

Then the imaginary equation is -2i=-iy?

OpenStudy (littlenugget):

yep, i think?

OpenStudy (anonymous):

to be sure, delete in the given equation all the terms that are not multiplied by "i", dont you get: -2i=-iy

OpenStudy (littlenugget):

ohhhh now i see

OpenStudy (anonymous):

Then, that means y=2, isnt it true?

OpenStudy (littlenugget):

is it because you divide both sides by i?

OpenStudy (anonymous):

Exactly! Now delete from the given equation all the terms multiplied by "i"

OpenStudy (littlenugget):

so it'd be 5-7=x-3

OpenStudy (anonymous):

Awesome! And what is the solution for the real part? x=???

OpenStudy (littlenugget):

1

OpenStudy (littlenugget):

i mean -1

OpenStudy (anonymous):

Do it again, what is the solution of 5-7=x+3?

OpenStudy (anonymous):

sorry 5-7=x-3

OpenStudy (littlenugget):

-2=x-3 +3 +3 ----------- x=1

OpenStudy (anonymous):

x=1, exactly! Have you got the rationale? Whenever you have an equation with complex numbers, obtain two equations, one for the real part and another for the imaginary part. And you know how to identify them already!

OpenStudy (anonymous):

don't you?

OpenStudy (littlenugget):

so basically for one of these problems, i just do what you told me?

OpenStudy (anonymous):

yes. Obtain the real part equation eliminating in the given equation all the terms multiplied by "i" and solve. And eliminate all the terms not multiplied by "i" in the given equation and solve.

OpenStudy (littlenugget):

ohhhhhhhhhhhh that makes so much more sense now!!!!! THANK YOU SO MUCH!!!!!! :DDDD

OpenStudy (anonymous):

you are more than welcome! good luck!